MILLIMETER-WAVE SPECTROSCOPY OF THE D_{2}CCD RADICAL

\underline{MITSUHIKO OHTSUKI}, MASATO HAYASHI, KENSUKE HARADA, KEIICHI TANAKA, Department of Chemistry, Faculty of Science, Kyushu University, Hakozaki, Higashiku, Fukuoka 812-8581, Japan.

The tunneling-rotation transitions of the vinyl-d_{3}(D_{2}CCD) radical produced by the 193 nm excimer laser photolysis of vinyl-d_{3} chloride(D_{2}CCDCl) have been observed by millimeter-wave spectroscopy combined with a pulsed supersonic jet technique. The b-type transitions, N_{K_{a}K_{c}}=1_{11}- 0_{00}, 1_{10}-1_{01}, 2_{12}-1_{01}, for both the 0^{−} ← 0^{+} and 0^{+} ← 0^{−} tunneling subbands were observed in the frequency region of 101-183 GHz, split into fine and hyperfine components due to the spin-rotation interaction and the spin-spin interaction of the α(CD)- and the β(CD_{2})-deuterons. The molecular constants such as rotational constants, spin-rotation interaction constants, and hyperfine interaction constants were determined by a simultaneous analysis of the observed spectrum and previously reported J=1←0 pure rotational transition^{a} together with the tunneling splitting ΔE_{0} = 771.843(23) MHz between the 0^{+} and 0^{−} states. The tunneling splitting for D_{2}CCD is less than 1/20 of that for H_{2}CCH (16 271.842 9(59) MHz)^{b} due to the mass effect of the α-H and D, and it is about 2/3 of H_{2}CCD (1 186.820(26) MHz) indicating the mixing of the vibrational modes for α-H/D and β-H_{2}/D_{2}. From the observed tunneling splitting, the barrier height h of the double minimum potential for D_{2}CCD was estimated to be 1549 cm^{−1} using one dimensional model. The barrier height h for D_{2}CCD is almost the same as those for H_{2}CCH and H_{2}CCD, 1580 and 1520 cm^{−1}, respectively, as expected by B.O. approximation and the isotopic effect due to zero point energies. The off-diagonal Fermi interaction constant, δa_{F}, which is responsible to the mixing of the wavefunctions of ortho(I_{β} = 0, 2) and para(I_{β} = 1) states, has been determined to be 19.8(30) MHz. The off-diagonal Fermi interaction may cause the nuclear spin conversion between the ortho- and para-states for D_{2}CCD.

aE. Kim, S. Yamamoto, J. Chem. Phys. \underline{116}, 10713, (2002).

bK. Tanaka, M. Toshimitsu, K. Harada, T. Tanaka, J. Chem. Phys. \underline{120}, 3604, (2004).